OpenStudy (anonymous):

10 - x -9= -11x + 91 + 4(3-x) please show how to solve

OpenStudy (anonymous):

10 - x -9= -11x + 91 + 4(3-x) 10 - x -9= -11x + 91 + 12 - 4x 10 - 9 - 91 - 12 = -11 x - 4x + x -102 = -14x x = 102 / 14

OpenStudy (anonymous):

answer will be a mixed #

OpenStudy (anonymous):

102 / 14 = 7 2/7

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

how did you get the 4x?

OpenStudy (anonymous):

4 (3-x) = 4*3 - 4*x = 12 - 4x